OpenStudy (anonymous):

solve. y2 − y + 5 = 0

OpenStudy (lgbasallote):

hint: use quadratic formula \[x = \frac{-b \pm \sqrt{b^2 - 4ac}}{2a}\] does that help?

OpenStudy (anonymous):

nah not really

OpenStudy (lgbasallote):

do you know what the quadratic formula is?

OpenStudy (anonymous):

yeah my teacher told us but he just gave us the homework and expected us to do this and turn it in tomorrow

OpenStudy (lgbasallote):

do you know how to use the quadratic formula?

OpenStudy (anonymous):

nope my first time

OpenStudy (lgbasallote):

but you have some basic background on it right? like the quadratic equation is in the form \[ax^2 + bx + c = 0\]

OpenStudy (anonymous):

somewhat lol

OpenStudy (lgbasallote):

okay for example you have \[x^2 + 2x + 3 = 0\] in this case a is 1 because the coefficient of x^2 here is 1 b is 2 because the coefficient of x us 2 and c is 3 because the constant is 3 so if i substitute a = 1; b = 2; c = 3 into \[x = \frac{-b \pm\sqrt{b^2 - 4ac}}{2a}\] \[ \implies x = \frac{(-2) \pm \sqrt{(2)^2 - 4(3)(1)}}{2(1)}\] \[\implies x = \frac{-2 \pm \sqrt{4 - 12}}{2}\] \[\implies x = \frac{-2 \pm \sqrt{-8}}{2}\] does that help?

OpenStudy (anonymous):

Somewhat... I'll try this out and see if I get the answer

OpenStudy (lgbasallote):

go for it

OpenStudy (anonymous):

go to mathway.com and type problem in

OpenStudy (anonymous):

it just shows the answer I need the step by step

OpenStudy (lgbasallote):

have you tried using the quadratic formula yet?

OpenStudy (anonymous):

yes sir, it's not working for me smh

OpenStudy (anonymous):

Math is my weakest subject

OpenStudy (lgbasallote):

lol i know your pain =))

OpenStudy (lgbasallote):

okay here's a hint.. in y^2 - y + 5 a = 1 b = -1 c = 5 does that help?

OpenStudy (anonymous):

so I'm writing y as a factor

OpenStudy (precal):

 if you look at the graph of your quadratic equation, you will notice that the function does not cross the x axis at all. That means that you have a complex solution

OpenStudy (precal):

you will not be able to factor this one at all

OpenStudy (anonymous):

now I see why I hate math

OpenStudy (precal):